# Aspects of the bulk flat space limit in AdS/CFT

### v1

@@ I Introduction
The AdS/CFT correspondence [1] is a remarkable duality between quantum field theories and quantum gravity. Basically stated, the dual CFT can be used as a definition of quantum gravity, in asymptotically AdS spaces. It is incumbent on us to find how geometry and locality in higher dimensions arises from the CFT dynamics. This is where the conjectures on the gap of anomalous dimensions arise from [2]. These can be used to prove from CFT assumptions that there is an effective dual effective theory in the bulk [3].
A natural notion of locality is that physics on scales much shorter than the AdS radius should be given by the physics of quantum gravity (or string theory) in flat space. It is understood that this description should be like the scattering S-matrix in flat space [4, 5]. An S-matrix should be thought of as a real time process in a Lorentzian geometry. The procedure of getting to this natural statement that arises from the equivalence principle is called the flat space limit. Our goal in this paper is to explain some of the kinematics of this limit in the bulk in a simpler way than what has been done before.
The standard way to approach calculations in AdS space and their dual CFT correlators utilizes an Euclidean setup with the GKPW dictionary [6, 7]. This obscures the Lorentzian physics required to understand small regions of spacetime, as well as the notion of in and out states. Many times these calculations are then analytically continued to try to get to Lorentzian physics.
A way to state the problem is as follows: what is the correct preparation of states to get the locally flat physics? At least in some setups this is considered problematic [8]. The other way to set up the problem is to begin this study by constructing a notion of a local bulk field. This is built by some integral representation on the boundary: the HKLL construction [9] and their earlier avatars [10]. This procedure is problematic when interactions are included, but it is considered as a way to address the issue.
@@
The main observation in our paper is that the flat space limit needs the following ingredients. A point $p$ in the bulk around which one will take a local flat space limit. The observation that the local Poincaré group at the point $p$ is a particular Ìnönü-Wigner contraction [11] of the AdS symmetry group. Because of this property, we can take the limit at the level of the representation theory of the AdS symmetry group. In that sense, we construct global wave functions for incoming and outgoing particles for the limit, rather than strictly local physics.
There is a technicality about representation theory of the conformal group (the AdS symmetry group) and its relation to the flat space limit of massive particles that is essential for our work. For massive states in the flat space limit, the concept of primary states of AdS survives, becoming a massive particle at rest. On the other hand, the descendants become null in the correct double scaling limit that leads to the Poincaré group. To recover the correct representation theory we have to work with primaries and their boosted states. These boosted states should be interpreted as a type of coherent state which becomes orthogonal to the primary in the flat space limit. The boosted states we construct end up providing exactly the particles at fixed momentum in the flat space limit and because of peculiarities of how they are constructed we can understand readily how they are prepared on the boundary.
@@
These states can then be easily assembled into integral representations of scattering amplitudes that are very suggestive of relations to other formulations of the flat space limit, not just in the strict AdS/CFT, but also for a (gapped) QFT in the flat space limit of curved AdS. An example of such works can be found in [12, 13, 14] (see also the Snowmass white paper [15]).
@@ II The Ìnönü-Wigner contraction and the flat space limit
The Poincaré group is one of the possible Ìnönü-Wigner contractions of the Conformal group of symmetries in AdS space, in the same sense than the group rigid translations of the plane is the Ìnönü-Wigner contraction of the group of symmetries of a sphere  [^1] Similarly, the Galilei group and the Carrollian group are Ìnönü-Wigner contractions of the Poincaré group [16].. Our goal is to explain in what sense this limit can be taken at the level of the representation theory of the group. For that, we need two elements: the concept of the Ìnönü-Wigner contraction as a limit and the proper representation theory of the group that we take a limit on. In this section we deal with the first : the group theory limit (the Lie theory of the appropriate vector fields in AdS).
@@ III The embedding space formalism
We will consider the flat space limit in the embedding space formalism. We are interested in the Lorentzian version of this problem, rather than the Euclidean version which has been developed in [17, 18].
@@ III.1 Scalar wave functions
The wave function on the left is implicitly living in ${{\mathbb{R}}}^{2,d}$, whereas the one on the right is strictly in AdS. They coincide on the hyperboloid. As a check, one can easily check that the wave function (21) solves the radial bulk wave equation
@@
arising from an AdS scalar field of mass squared $m^{2}=\Delta(\Delta-d)$ and frequency $\Delta$. Notice $G(\rho)=\cosh\rho\,(\sinh\rho)^{d-1}$ is the AdS volume form $\sqrt{-\det g}$. Mathematically, the wave function $z^{-\Delta}$ is holomorphic (harmonic) outside the disk $|z|\geq 1$. It actually lives in the universal cover of the AdS hyperboloid as $\Delta$ is a real number. It would have been multi-valued otherwise. Notice how this construction simplifies the analysis relative to other works (for example [19], which uses a different coordinate system).
@@
Thus, the wave function becomes negligible when $\rho\gg 1/\sqrt{\Delta}$. In our units, a distance of order one in $\rho$ is order one in the radius of AdS. Hence for large $\Delta$, the wave function is localized on distances much shorter than the AdS scale. Indeed, if $\Delta$ is large, one can think of a particle at the bottom of AdS as a particle in a harmonic oscillator trap [20], whose ground state wave function is Gaussian.
@@
Primary wave functions of large $\Delta$ can be thought of as (semi) classical particles of mass $m_{AdS}$ in the minimal energy allowed: a particle at rest at $\rho=0$. In such a limit, the energy of the particle $\Delta$ is roughly $m_{AdS}+d/2$. The factor of $d/2$ can be interpreted as a zero point energy from thinking of AdS as a harmonic trap (see for example [21]).
@@
where $f_{s}$ is a polynomial of degree $s$ in the combined variables. The $\psi_{\text{desc}}$ are (generalized) harmonic functions in the embedding space ${\mathbb{R}}^{2,d}$ and carry energy $E=\Delta+s$. This is a generalization to AdS of the construction of spherical harmonics as harmonic polynomials in an embedding space (see the discussion in [22] and references therein, which we found useful). We can then restrict to the AdS hyperboloid and rewrite these in terms of the $\rho,t,\Omega$ variables if we want to.
@@ IV The flat space limit in the embedding space formalism and representation theory
Once the wave functions for a particle at rest are known, the wave functions for a moving particle are generated by boosting them (this is part of the standard foundations in QFT and it is especially well presented in [23]). This procedure amounts to a simple change of coordinates by a Lorentz transformation $(x^{\mu})^{\prime}=\Lambda^{\mu}_{\nu}x^{\nu}$. Once more, the embedding space formalism described earlier allows us to perform this boost globally in AdS, and not just locally at $p$. Also, since the action involves an AdS isometry, it preserves norms, i.e. it is a unitary transformation in the full Hilbert space of single particle states in AdS. This last point is crucial, as once the norm of the primary state is fixed, all other states obtained by rotation this way have a known norm. Via the AdS/CFT correspondence, this must be the same as the norm of the state in the dual CFT.
@@
Some readers may recognise these wave forms as an analytic continuation of wave functions in Euclidean AdS [17], where the notion of $\kappa$ has also been analytically continued.
@@
where we used $y^{-1}=\frac{1}{2}(z+\bar{z}))$ and $y^{0}=\frac{1}{2i}(z-\bar{z})$. Since $|z|>1$, the boosted AdS wave function can be Taylor expanded in powers of $1/z$ as follows
@@ V Amplitudes and state preparation
The idea of an amplitude is to set it up in time dependent perturbation theory for quantum field in curved space assuming that it is like a finite box. Since AdS has a global timelike Killing vector, we can canonically decompose the states into positive and negative energy and some of the usual complications of doing quantum field theory in curved space are absent (see the book by Birrell and Davis [24]). We thus write
@@
where $\lambda$ is the perturbative coupling constant associated to the potential. In essence, we are doing Feynman diagrams in the bulk with wave functions for external particles at finite time. Within the LSZ formalism, one usually works with Green’s functions and amputates the external legs replacing them by wave functions. Since the current wave functions are already known, there is no need to write the LSZ integrals defining them (see for example [25], chapter 5 and [26], chapter 6).
@@
This will occur also to higher orders in perturbation theory. So long as the corresponding Feynman propagators are sufficiently suppressed at long distances, the amplitudes can be just as well computed in global AdS as in flat space.
@@
The turning points are important because they are naturally associated with Euclidean preparation of the corresponding one particle state. The idea is very simple. At the turning point all spatial momenta vanish, so it is easy to set up an analytic continuation $i\vec{p}_{E}=\vec{p}_{L}$ that keeps $p$ real both for the Euclidean and the Lorentzian theory, like is usually done in a Hartle-Hawking wave function (see also the appendix in [27] to undestand how to deal with angular momentum).
@@
As was computed in [28], we find that the Euclidean time in the boundary (towards the past) and the radius of the turning point $\rho^{*}$ are related by
@@
so that there is a precise relation between the geometry of $\rho^{*}$ and the boundary time insertion, which is independent of $\Delta$. Given that the energy (in AdS units) of the particle at $\rho^{*}$ is $E=\Delta\cosh(\rho^{*})=\Delta\cosh(\eta)$ where $\eta$ is the rapidity at the origin, we find that knowledge of $\rho^{*}$ is equivalent to the knowledge of the rapidity in the flat space limit. Similarly, the value of the Euclidean time $\tau$ on the boundary completely determines $\rho^{*}$. The angles on the sphere are also the same, so the Euclidean time position plus the angle completely specifies the incoming or outgoing states. A similar picture arises also in the work [29] where the in/out states are built from boundary Euclidean caps on AdS and the boundary to bulk Green’s functions take a similar form to ours.
@@
Therefore the correct prescription is to include a factor of $(2\sinh(\tau))^{\Delta}$ in front of the operator insertion. This makes sure that we have properly normalized the states in the boundary at $t=-\frac{\pi}{2}$ to correspond to the same normalization as states in the bulk (a similar statement is needed at $t=\frac{\pi}{2}$ for the out states). This factor $(2\sinh(\tau))^{-\Delta}$ is the exponential of the subtracted action for the spacelike geodesic that prepares the state [28]. It is essentially $\exp(-S)$. It can be thought of as the propagator evaluated from the boundary to the turning point.
@@ VI Massless particles and their wave functions
The first one is associated to the hard preparation of scattering states [8]. In principle, constructions like HKLL [9] can provide the local fields, and these can be convoluted with the desired wave function to prepare the states. This is the approach of Hijano [30]. In a sense, we would be doing the reverse of (8).
@@
Notice that $(1-iv_{\mu}x^{\mu})^{-\Delta}$ is a non-singular solution of the wave equation in the limit. The wave function has singularities at complex values of $x^{\mu}$, but it is perfectly regular in flat space . Also, for fixed $\Delta$, remember that the wave functions have a width of order one in $AdS$. The large boost has Lorentz contracted this distance to be finite in the longitudinal direction of the wave profile fitting nicely inside the flat space region near the point $p$. These can be considered very sharp waves in the direction of $\hat{n}$, but they are wide functions in the transverse directions to the front of the wave. These wave functions are more similar to the ones that arise when considering celestial amplitudes (see for example the reviews [31, 32]), although they are not the exact same ones. For example, in our case $\Delta$ is real and positive, whereas in the celestial wave functions one usually has a parameter $\Delta$ that is complex and the one in the denominator is missing. Notice that $\Delta$ instead of becoming a mass parameter, appears in the limit as part of the shape of the wave packets. Different values of $\Delta$ are still associated to different species of massless particles since they are considered different species in AdS after all.
@@
Notice that since $v_{\nu}$ is null, the integrand itself is a solution of the massless Klein Gordon equation. These look like a Mellin integral, so we will call them such. The expressions then have a resemblance to the Mellin amplitude formalism of Penedones et al. [12, 33].
We can now try and compute amplitudes as we did in section V. We just need to replace the integrals for time dependent perturbation theory, using the wave packets we have for in and out states. Amplitudes involving the $\psi$ can then be rewritten as integrals over the S-matrix, with one Schwinger parameter per external particle. Indeed, if we write the in and out states in terms of waves at fixed momentum with the aid of (58), this statement is self-evident. We will get this way a generalized notion of Fourier transform of the S-matrix (this will be different than the celestial transform [34]).
We want to argue that only a small region near the origin contributes. We will take that as a given in our formulae, but this is actually very subtle. The reason that it works here, unlike the case for very massive particles, is that the wave profiles decay for large spatial $x$. Two incoming wave packets that collide will both have large values of the field in some spatial hypersurface, but this does not extend in time. Hence the time interval where the interactions happen is of small duration, even in the $x$ coordinates. The spatial region still goes to infinity and can extend to AdS size radii transversely. If enough of the outgoing particle wave functions intersect transversely to each other in this hyperplane, then we should be able to localize the multi-particle amplitude to a small region of flat space, finding an approximation of Landau singularities in the full AdS as in [35] (alternatively, lightcone cuts for measurements intersecting in the bulk at the point $p$ as in [36]).
@@
depending on the Schwinger parameters. Notice the resulting massless amplitude (59) is a specific Mellin integral of $\exp(-\sum s_{i})\delta^{d+1}(\sum s_{i}v_{\mu})S_{\text{massless}}(sv_{\mu,i})$ (a Mellin transform evaluated at specific values determined by the set of $\Delta$’s). For generic null momentum vectors $v_{\nu}$, with a small number of particles, the delta functions eliminate most, but not all of the $s$ variables. Since given a solution of the delta function support, one can obtain a different one by rescaling all the vectors by the same amount $s_{i}\to\gamma s_{i}$, we infer that after evaluating the delta function in (59), there should still be at least one Mellin integral left to perform. If one of the particles is massive, the scale can (at least in principle) may be resolved by the massive particle momentum and its recoil.
@@
$\tau$ is infinitesimally small (of order $1/\Lambda$ in our notation). Hence, the flat space limit for massless particles on the boundary requires a double scaling limit of the insertion point of the operator, where the Euclidean time evolution goes to zero with the same speed as the flat space limit parameter $1/\Lambda$. In other words, the insertions of the corresponding operators occur almost in real time. The scattering events then look very similar to state preparation that occurs in a small band around $t=-\frac{\pi}{2}$ and readout at $t=\frac{\pi}{2}$. If we were to prepare the state à la HKLL, the insertion can be reduced to a very small band in space and time around the time where match to Euclidean physics. This is the same geometric limit as in [37] and corresponds to Landau singularities in the bulk [35]. Similar limits appear in constructions like the ones in [38], which addresses some of the complications of thinking about the flat space S-matrix.
A more recent expression of conjectures related to the flat space limit in the language of conformal correlators can be found in [39]. These seem to depend on being able to work with complex momenta. Our description works for any $\Delta$, but the precise double scaling on the boundary that we found that produces (56) is new. Another approach starts from CFT and extracts unitarity of the S-matrix from conformal correlators [40]. Again, these all are correlators at finite $\Delta$, so all particles become massless in such a limit and are subject to our geometric description in the bulk.

### v2

@@ I Introduction
The AdS/CFT correspondence [Maldacena:1997re] is a remarkable duality between quantum field theories and quantum gravity. Basically stated, the dual CFT can be used as a definition of quantum gravity, in asymptotically AdS spaces. It is incumbent on us to find how geometry and locality in higher dimensions arises from the CFT dynamics. This is where the conjectures on the gap of anomalous dimensions arise from [Heemskerk:2009pn]. These can be used to prove from CFT assumptions that there is an effective dual effective theory in the bulk [Caron-Huot:2021enk].
A natural notion of locality is that physics on scales much shorter than the AdS radius should be given by the physics of quantum gravity (or string theory) in flat space. It is understood that this description should be like the scattering S-matrix in flat space [Polchinski:1999ry, Susskind:1998vk]. An S-matrix should be thought of as a real time process in a Lorentzian geometry. The procedure of getting to this natural statement that arises from the equivalence principle is called the flat space limit. Our goal in this paper is to explain some of the kinematics of this limit in the bulk in a simpler way than what has been done before.
The standard way to approach calculations in AdS space and their dual CFT correlators utilizes an Euclidean setup with the GKPW dictionary [Gubser:1998bc, Witten:1998qj]. This obscures the Lorentzian physics required to understand small regions of spacetime, as well as the notion of in and out states. Many times these calculations are then analytically continued to try to get to Lorentzian physics. For a recent example see [Marotta:2024sce], where a 3 vector interaction vertex is treated in detail and it is argued how the flat space S-matrix arises from an analytic continuation of the Euclidean answer that they compute.
A way to state the problem is as follows: what is the correct preparation of states to get the locally flat physics? At least in some setups this is considered problematic [Giddings:1999jq]. The other way to set up the problem is to begin this study by constructing a notion of a local bulk field. This is built by some integral representation on the boundary: the HKLL construction [Hamilton:2006az] and their earlier avatars [Banks:1998dd]. This procedure is problematic when interactions are included, but it is considered as a way to address the issue.
@@
The main observation in our paper is that the flat space limit needs the following ingredients. A point $p$ in the bulk around which one will take a local flat space limit. The observation that the local Poincaré group at the point $p$ is a particular Ìnönü-Wigner contraction [Inonu:1953sp] of the AdS symmetry group. Because of this property, we can take the limit at the level of the representation theory of the AdS symmetry group. In that sense, we construct global wave functions for incoming and outgoing particles for the limit, rather than strictly local physics.
There is a technicality about representation theory of the conformal group (the AdS symmetry group) and its relation to the flat space limit of massive particles that is essential for our work. For massive states in the flat space limit, the concept of primary states of AdS survives, becoming a massive particle at rest. On the other hand, the descendants become null in the correct double scaling limit that leads to the Poincaré group. To recover the correct representation theory we have to work with primaries and their boosted states. These boosted states should be interpreted as a type of coherent state which becomes orthogonal to the primary in the flat space limit. The boosted states we construct end up providing exactly the particles at fixed momentum in the flat space limit and because of peculiarities of how they are constructed we can understand readily how they are prepared on the boundary [^1] There is a broad literature studying the representation theory of the conformal group, the Poincaré group and its carrollian avatars. We would like to point out [Iacobacci:2024laa], where the authors explain how to decompose massless and massive particles into unitary irreducible representations of the Lorentz group, for arbitrary spin, yielding the Mellin transform that appears in celestial holography. Also, [Nguyen:2023vfz] builds conformal field representations of the Poincaré group living at null infinity (à la Mack-Salam) and shows how this relates to massless particles via Fourier-Mellin transforms..
@@
These states can then be easily assembled into integral representations of scattering amplitudes that are very suggestive of relations to other formulations of the flat space limit, not just in the strict AdS/CFT, but also for a (gapped) QFT in the flat space limit of curved AdS. An example of such works can be found in [Penedones:2010ue, Paulos:2016fap, vanRees:2022zmr] (see also the Snowmass white paper [Kruczenski:2022lot]).
@@
Before finishing this introduction, we would like to explicitly mention the existence of a large literature discussing the flat limit of the AdS/CFT correspondence, both from the bulk and from the boundary perspectives. For a review on the carrollian structure emerging in the boundary theory, we refer the readers to [Bagchi:2025vri]. For reviews on celestial holography, see [Raclariu:2021zjz, Pasterski:2021rjz]. For a non-exhaustive list of references discussing bulk aspects of the limit, as in this manuscript, see [Fitzpatrick:2010zm, Compere:2019bua, Compere:2020lrt, Campoleoni:2023fug, Bekaert:2024itn, Marotta:2024sce, Alday:2024yyj, Lipstein:2025jfj] and references therein.
@@ II The Ìnönü-Wigner contraction and the flat space limit
The Poincaré group is one of the possible Ìnönü-Wigner contractions of the Conformal group of symmetries in AdS space, in the same sense than the group rigid translations of the plane is the Ìnönü-Wigner contraction of the group of symmetries of a sphere  [^2] Similarly, the Galilei group and the Carrollian group are Ìnönü-Wigner contractions of the Poincaré group [levy1965nouvelle].. Our goal is to explain in what sense this limit can be taken at the level of the representation theory of the group. For that, we need two elements: the concept of the Ìnönü-Wigner contraction as a limit and the proper representation theory of the group that we take a limit on. In this section we deal with the first : the group theory limit (the Lie theory of the appropriate vector fields in AdS).
@@ III The embedding space formalism
We will consider the flat space limit in the embedding space formalism. We are interested in the Lorentzian version of this problem, rather than the Euclidean version which has been developed in [Costa:2011mg, Costa:2014kfa].
@@ III.1 Scalar wave functions
The wave function on the left is implicitly living in ${{\mathbb{R}}}^{2,d}$, whereas the one on the right is strictly in AdS. They coincide on the hyperboloid. As a check, one can easily verify that the wave function (21) solves the radial bulk wave equation
@@
arising from an AdS scalar field of mass squared $m^{2}=\Delta(\Delta-d)$ and frequency $\Delta$. Notice $G(\rho)=\cosh\rho\,(\sinh\rho)^{d-1}$ is the AdS volume form $\sqrt{-\det g}$. Mathematically, the wave function $z^{-\Delta}$ is holomorphic (harmonic) outside the disk $|z|\geq 1$. It actually lives in the universal cover of the AdS hyperboloid as $\Delta$ is a real number. It would have been multi-valued otherwise. Notice how this construction simplifies the analysis relative to other works (for example [Fitzpatrick:2010zm], which uses a different coordinate system).
@@
Thus, the wave function becomes negligible when $\rho\gg 1/\sqrt{\Delta}$. In our units, a distance of order one in $\rho$ is order one in the radius of AdS. Hence for large $\Delta$, the wave function is localized on distances much shorter than the AdS scale. Indeed, if $\Delta$ is large, one can think of a particle at the bottom of AdS as a particle in a harmonic oscillator trap [Berenstein:2002jq], whose ground state wave function is Gaussian.
@@
Primary wave functions of large $\Delta$ can be thought of as (semi) classical particles of mass $m_{AdS}$ in the minimal energy allowed: a particle at rest at $\rho=0$. In such a limit, the energy of the particle $\Delta$ is roughly $m_{AdS}+d/2$. The factor of $d/2$ can be interpreted as a zero point energy from thinking of AdS as a harmonic trap (see for example [Berenstein:2002ke]).
@@
where $f_{s}$ is a polynomial of degree $s$ in the combined variables. The $\psi_{\text{desc}}$ are (generalized) harmonic functions in the embedding space ${\mathbb{R}}^{2,d}$ and carry energy $E=\Delta+s$. This is a generalization to AdS of the construction of spherical harmonics as harmonic polynomials in an embedding space (see the discussion in [vanNieuwenhuizen:2012zk] and references therein, which we found useful). We can then restrict to the AdS hyperboloid and rewrite these in terms of the $\rho,t,\Omega$ variables if we want to.
@@ IV The flat space limit in the embedding space formalism and representation theory
Once the wave functions for a particle at rest are known, the wave functions for a moving particle are generated by boosting them (this is part of the standard foundations in QFT and it is especially well presented in [Weinberg:1995mt]). This procedure amounts to a simple change of coordinates by a Lorentz transformation $(x^{\mu})^{\prime}=\Lambda^{\mu}_{\nu}x^{\nu}$. Once more, the embedding space formalism described earlier allows us to perform this boost globally in AdS, and not just locally at $p$. Also, since the action involves an AdS isometry, it preserves norms, i.e. it is a unitary transformation in the full Hilbert space of single particle states in AdS. This last point is crucial, as once the norm of the primary state is fixed, all other states obtained by rotation this way have a known norm. Via the AdS/CFT correspondence, this must be the same as the norm of the state in the dual CFT.
@@
Some readers may recognise these wave forms as an analytic continuation of wave functions in Euclidean AdS [Costa:2011mg], where the notion of $\kappa$ has also been analytically continued.
@@
where we used $y^{-1}=\frac{1}{2}(z+\bar{z})$ and $y^{0}=\frac{1}{2i}(z-\bar{z})$. Since $|z|>1$, the boosted AdS wave function can be Taylor expanded in powers of $1/z$ as follows
@@ V Amplitudes and state preparation
The idea of an amplitude is to set it up in time dependent perturbation theory for quantum field in curved space assuming that it is like a finite box. Since AdS has a global timelike Killing vector, we can canonically decompose the states into positive and negative energy and some of the usual complications of doing quantum field theory in curved space are absent (see the book by Birrell and Davis [Birrell:1982ix]). We thus write
@@
where $\lambda$ is the perturbative coupling constant associated to the potential. In essence, we are doing Feynman diagrams in the bulk with wave functions for external particles at finite time. Within the LSZ formalism, one usually works with Green’s functions and amputates the external legs replacing them by wave functions. Since the current wave functions are already known, there is no need to write the LSZ integrals defining them (see for example [Srednicki:2007qs], chapter 5 and [Schwartz:2014sze], chapter 6).
@@
This will occur also to higher orders in perturbation theory. So long as the corresponding Feynman propagators are sufficiently suppressed at long distances, the amplitudes can be just as well computed in global AdS as in flat space. This will be true Feynman diagram by diagram  [^5] We should note that there are various claims in the literature about the flat limit matching Feynman diagram by Feynman diagram, starting from Euclidean correlators and performing analytic continuations, for example [Fitzpatrick:2011hu, Marotta:2024sce]. In our opinion, a common feature of these works is that the real time aspect that only a small region of space and time contributes is obscured..
@@
The turning points are important because they are naturally associated with Euclidean preparation of the corresponding one particle state. The idea is very simple. At the turning point all spatial momenta vanish, so it is easy to set up an analytic continuation $i\vec{p}_{E}=\vec{p}_{L}$ that keeps $p$ real both for the Euclidean and the Lorentzian theory, like is usually done in a Hartle-Hawking wave function (see also the appendix in [Berenstein:2020vlp] to understand how to deal with angular momentum).
@@
As was computed in [Berenstein:2019tcs], we find that the Euclidean time in the boundary (towards the past) and the radius of the turning point $\rho^{*}$ are related by
@@
so that there is a precise relation between the geometry of $\rho^{*}$ and the boundary time insertion, which is independent of $\Delta$. Given that the energy (in AdS units) of the particle at $\rho^{*}$ is $E=\Delta\cosh(\rho^{*})=\Delta\cosh(\eta)$ where $\eta$ is the rapidity at the origin, we find that knowledge of $\rho^{*}$ is equivalent to the knowledge of the rapidity in the flat space limit. Similarly, the value of the Euclidean time $\tau$ on the boundary completely determines $\rho^{*}$. The angles on the sphere are also the same, so the Euclidean time position plus the angle completely specifies the incoming or outgoing states. A similar picture arises also in the work [Komatsu:2020sag] where the in/out states are built from boundary Euclidean caps on AdS and the boundary to bulk Green’s functions take a similar form to ours.
@@
Therefore the correct prescription is to include a factor of $(2\sinh(\tau))^{\Delta}$ in front of the operator insertion. This makes sure that we have properly normalized the states in the boundary at $t=-\frac{\pi}{2}$ to correspond to the same normalization as states in the bulk (a similar statement is needed at $t=\frac{\pi}{2}$ for the out states). This factor $(2\sinh(\tau))^{-\Delta}$ is the exponential of the subtracted action for the spacelike geodesic that prepares the state [Berenstein:2019tcs]. It is essentially $\exp(-S)$. It can be thought of as the propagator evaluated from the boundary to the turning point.
@@ VI Massless particles and their wave functions
The first one is associated to the hard preparation of scattering states [Giddings:1999jq]. In principle, constructions like HKLL [Hamilton:2006az] can provide the local fields, and these can be convoluted with the desired wave function to prepare the states. This is the approach of Hijano [Hijano:2019qmi]. In a sense, we would be doing the reverse of (8).
@@
Notice that $(1-iv_{\mu}x^{\mu})^{-\Delta}$ is a non-singular solution of the wave equation in the limit. The wave function has singularities at complex values of $x^{\mu}$, but it is perfectly regular in flat space. Also, for fixed $\Delta$, remember that the wave functions have a width of order one in $AdS$. The large boost has Lorentz contracted this distance to be finite in the longitudinal direction of the wave profile fitting nicely inside the flat space region near the point $p$. These can be considered very sharp waves in the direction of $\hat{n}$, but they are wide functions in the transverse directions to the front of the wave. These wave functions are more similar to the ones that arise when considering celestial amplitudes (see for example the reviews [Raclariu:2021zjz, Pasterski:2021rjz]), although they are not the exact same ones. For example, in our case $\Delta$ is real and positive, whereas in the celestial wave functions one usually has a parameter $\Delta$ that is complex and the one in the denominator is missing. Notice that $\Delta$ instead of becoming a mass parameter, appears in the limit as part of the shape of the wave packets. Different values of $\Delta$ are still associated to different species of massless particles since they are considered different species in AdS after all.
@@
Notice that since $v_{\nu}$ is null, the integrand itself is a solution of the massless Klein Gordon equation. These look like a Mellin integral, so we will call them such. The expressions then have a resemblance to the Mellin amplitude formalism of Penedones et al. [Penedones:2010ue, Fitzpatrick:2011ia].
We can now try and compute amplitudes as we did in section V. We just need to replace the integrals for time dependent perturbation theory, using the wave packets we have for in and out states. Amplitudes involving the $\psi$ can then be rewritten as integrals over the S-matrix, with one Schwinger parameter per external particle. Indeed, if we write the in and out states in terms of waves at fixed momentum with the aid of (58), this statement is self-evident. We will get this way a generalized notion of Fourier transform of the S-matrix (this will be different than the celestial transform [Pasterski:2016qvg]).
We want to argue that only a small region near the origin contributes. We will take that as a given in our formulae, but this is actually very subtle. The reason that it works here, unlike the case for very massive particles, is that the wave profiles decay for large spatial $x$. Two incoming wave packets that collide will both have large values of the field in some spatial hypersurface, but this does not extend in time. Hence the time interval where the interactions happen is of small duration, even in the $x$ coordinates. The spatial region still goes to infinity and can extend to AdS size radii transversely. If enough of the outgoing particle wave functions intersect transversely to each other in this hyperplane, then we should be able to localize the multi-particle amplitude to a small region of flat space, finding an approximation of Landau singularities in the full AdS as in [Maldacena:2015iua] (alternatively, lightcone cuts for measurements intersecting in the bulk at the point $p$ as in [Engelhardt:2016wgb]).
@@
depending on the Schwinger parameters. Notice the resulting massless amplitude (59) is a specific Mellin integral of $\exp(-\sum s_{i})\delta^{d+1}(\sum s_{i}v_{\mu})S_{\text{massless}}(sv_{\mu,i})$ (a Mellin transform evaluated at specific values determined by the set of $\Delta$’s). For generic null momentum vectors $v_{\nu}$, with a small number of particles, the delta functions eliminate most, but not all of the $s$ variables. Since given a solution of the delta function support, one can obtain a different one by rescaling all the vectors by the same amount $s_{i}\to\gamma s_{i}$, we infer that after evaluating the delta function in (59), there should still be at least one Mellin integral left to perform. If one of the particles is massive, the scale can (at least in principle) be resolved by the massive particle momentum and its recoil.
@@
$\tau$ is infinitesimally small (of order $1/\Lambda$ in our notation). Hence, the flat space limit for massless particles on the boundary requires a double scaling limit of the insertion point of the operator, where the Euclidean time evolution goes to zero with the same speed as the flat space limit parameter $1/\Lambda$. In other words, the insertions of the corresponding operators occur almost in real time. The scattering events then look very similar to state preparation that occurs in a small band around $t=-\frac{\pi}{2}$ and readout at $t=\frac{\pi}{2}$. If we were to prepare the state à la HKLL, the insertion can be reduced to a very small band in space and time around the time where we match to Euclidean physics. This is the same geometric limit as in [deGioia:2024yne] and corresponds to Landau singularities in the bulk [Maldacena:2015iua]. Similar limits appear in constructions like the ones in [Gary:2009mi], which addresses some of the complications of thinking about the flat space S-matrix.
A more recent expression of conjectures related to the flat space limit in the language of conformal correlators can be found in [vanRees:2023fcf]. These seem to depend on being able to work with complex momenta. Our description works for any $\Delta$, but the precise double scaling on the boundary that we found that produces (56) is new. Another approach starts from CFT and extracts unitarity of the S-matrix from conformal correlators [Fitzpatrick:2011dm]. Again, these all are correlators at finite $\Delta$, so all particles become massless in such a limit and are subject to our geometric description in the bulk.